In 8 hours, an experienced cook prepares enough pies to supply a local restaurant’s daily order. Another cook prepares the same number of pies in 9 hours. Together with a third cook, they prepare the pies in 3 hours. Find how long it takes the third cook to prepare the pies alone.
Answer by jorel1380(3719)

1/8+1/9+1/n=1/3
1/n=1/3-1/8-1/9
"=24/72-9/72-8/72=24-17/72=7/72
7n=72
n=72/7
It takes the third cook 10-2/7 hours to cook the pies alone..